OpenStudy (anonymous):

Simplify the number using imaginary unit i. √-16 ^that's square root -16. PLEASE help! I will give medal to best answer! Show all steps please!

OpenStudy (anonymous):

Square root of 16 is easy! 4 Square root of a negative number means you will get "i" So we will get 4i

OpenStudy (anonymous):

That's all the steps to the problem? Wow! I feel dumb for needing to post this, lol. Thank you!!

OpenStudy (jdoe0001):

\(\bf\color{red}{ i = \sqrt{-1}}\qquad thus\\ \quad \\ \sqrt{-16}\implies \sqrt{16}\times \sqrt{-1}\implies \sqrt{4^2}\times \sqrt{-1}\implies 4\times i\implies 4i\)

OpenStudy (anonymous):

Its okay! Don't feel dumb, @jdoe0001 has listed a really nice solution for you. Just split up the "i" portion of the problem and the sqrt portion of the problem and they will be a bit easier!

OpenStudy (anonymous):

Could one of you help me with another problem?

OpenStudy (anonymous):

Sure! Go ahead and message me your link

OpenStudy (anonymous):

Simplify the expression. -5+i/2i *that's in fraction form*

OpenStudy (anonymous):

@jdoe0001 @jim_thompson5910

OpenStudy (anonymous):

@mathstudent55 @mathman806

jimthompson5910 (jim_thompson5910):

Hint: multiply top and bottom by 'i' \[\large \frac{-5+i}{2i}\] \[\large \frac{(-5+i)*i}{2i*i}\] I'll let you finish up

OpenStudy (anonymous):

\[\frac{-5i^{2}}{2i^2 }\]

OpenStudy (anonymous):

Right @jim_thompson5910?

jimthompson5910 (jim_thompson5910):

close, the numerator should be \(\large -10i + 2i^2\)

jimthompson5910 (jim_thompson5910):

there are more steps after that hint: \(\large i^2 = -1\)

OpenStudy (anonymous):

\[\frac{ 5 }{ 8 }\]

jimthompson5910 (jim_thompson5910):

not sure how you got that

OpenStudy (anonymous):

i subbed in -1 for all i terms

OpenStudy (jdoe0001):

\(\bf \cfrac{(-5+i)\times i}{2i\times i}\implies \cfrac{-5i+i^2}{2i^2}\)

OpenStudy (jdoe0001):

\(\bf \cfrac{(-5+i)\times i}{2i\times i}\implies\cfrac{(-5\times i)+(i\times i)}{2i^2}\implies \cfrac{-5i+i^2}{2i^2}\)

OpenStudy (jdoe0001):

\(\bf \textit{recall that}\\ \quad \\ i^2\implies \sqrt{1}\times \sqrt{1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1\)

OpenStudy (jdoe0001):

ahemm... well.. rather \(\bf i^2\implies \sqrt{-1}\times \sqrt{-1}\implies (\sqrt{-1})^2\implies \sqrt{(-1)^2}\implies -1\)